# Travelling Waves in a Mathematical Model for Oncolytic Virotherapy

### v1

@@ 5 Numerics
In this section we show how our theoretical minimum wave speeds compare to the numerically observed wave speeds. To address this, in Figure 4, we present travelling–wave simulations of the model, carried out in Matlab using the pdepe solver. On the left column of Figure (4), we run the simulations using the parameter values used previously by [arwa]. In that case the parameter $a$ was greater than one and for $\gamma$ they used the value $\frac{40}{3}$. We compare the difference in wave formation for smaller $a$ value in the right column of Figure 4.
@@ 6 Conclusion
In this paper, we study a reaction–diffusion model designed to capture the spatiotemporal dynamics of oncolytic virotherapy (OVT) and tumor cell populations. The model, originally proposed by Baabdulla et al. [arwa], describes the interactions between uninfected cancer cells, infected cancer cells, and free virions. A central open question in OVT is whether the virus, when administered as a monotherapy, can spread through and eradicate the tumor before the tumor itself invades the surrounding tissue. Mathematical modelling provides a natural framework to explore this question by comparing the invasion speed of the virus with that of the cancer cells, and by identifying which viral parameters could be enhanced to tip the competition in favour of successful tumor clearance. In this work, we rigorously establish the existence of traveling–wave solutions for this non-cooperative three-component system, proving that such waves exist for all wave speeds $c\geq\bar{c}$ . While traveling waves have been extensively studied for both cooperative and non-cooperative systems, the present model poses additional challenges due to its three interacting variables and non-monotone structure, which impose technical restrictions in the construction of upper and lower solutions. As a result, the proof naturally leads to parameter regimes where existence can be established and others where the status of travelling waves remains unclear. These “foggy” regions highlight interesting mathematical questions for future investigation, particularly regarding whether the remaining gaps can be closed by refining the analytical framework. Finally, to complement the theoretical analysis, we conduct numerical simulations that illustrate the full PDE dynamics, confirm the formation of travelling invasion fronts, and demonstrate how wave propagation depends on the viral parameters. Together, the analytical and numerical results provide a deeper understanding of the mechanisms that influence the success of OVT as a monotherapy and point toward avenues for optimizing viral design.

### v2

@@ 5 Numerics
In this section we show how our theoretical minimum wave speeds compare to the numerically observed wave speeds. To address this, in Figure 4, we present travelling–wave simulations of the model, carried out in Matlab using the pdepe solver. On the left column of Figure (4), we run the simulations using the parameter values used previously by [2]. In that case the parameter $a$ was greater than one and for $\gamma$ they used the value $\frac{40}{3}$. We compare the difference in wave formation for smaller $a$ value in the right column of Figure 4.
@@ 6 Conclusion
In this paper, we study a reaction–diffusion model designed to capture the spatiotemporal dynamics of oncolytic virotherapy (OVT) and tumor cell populations. The model, originally proposed by Baabdulla et al. [2], describes the interactions between uninfected cancer cells, infected cancer cells, and free virions. A central open question in OVT is whether the virus, when administered as a monotherapy, can spread through and eradicate the tumor before the tumor itself invades the surrounding tissue. Mathematical modelling provides a natural framework to explore this question by comparing the invasion speed of the virus with that of the cancer cells, and by identifying which viral parameters could be enhanced to tip the competition in favour of successful tumor clearance. In this work, we rigorously establish the existence of traveling–wave solutions for this non-cooperative three-component system, proving that such waves exist for all wave speeds $c\geq\bar{c}$ . While traveling waves have been extensively studied for both cooperative and non-cooperative systems, the present model poses additional challenges due to its three interacting variables and non-monotone structure, which impose technical restrictions in the construction of upper and lower solutions. As a result, the proof naturally leads to parameter regimes where existence can be established and others where the status of travelling waves remains unclear. These “foggy” regions highlight interesting mathematical questions for future investigation, particularly regarding whether the remaining gaps can be closed by refining the analytical framework. Finally, to complement the theoretical analysis, we conduct numerical simulations that illustrate the full PDE dynamics, confirm the formation of travelling invasion fronts, and demonstrate how wave propagation depends on the viral parameters. Together, the analytical and numerical results provide a deeper understanding of the mechanisms that influence the success of OVT as a monotherapy and point toward avenues for optimizing viral design.